OpenStudy (jazzyfa30):

Answer maximum value of ____ or minimum value of (blank) Determine whether each function has a maximum or a minimum value. Then find the maximum or minimum value of each function. f(x) = x2 +9

OpenStudy (anonymous):

minimum value is 9. Since x^2 at most can be zero, not less.

OpenStudy (rsadhvika):

f(x) = ax^2 + bx + c

OpenStudy (rsadhvika):

a > 0 => MINIMUM value a < 0 => MAXIMUM value

OpenStudy (rsadhvika):

here a = 1, which is > 0, so it will have MINIMUM value

OpenStudy (jazzyfa30):

@rsadhvika can u walk me through it

OpenStudy (rsadhvika):

f(x) = x2 +9 compare it with ax2 + bx + c

OpenStudy (rsadhvika):

a=1, right ?

OpenStudy (jazzyfa30):

yes

OpenStudy (rsadhvika):

first we figure out whether it has MAXIMUM or MINIMUM value a > 0 => MINIMUM value a < 0 => MAXIMUM value

OpenStudy (rsadhvika):

since a = 1, its greater than 0, so it will have MINIMUM value. fine so far ?

OpenStudy (jazzyfa30):

yes

OpenStudy (rsadhvika):

Having known it has a MINIMUM value, lets find the MINIMUM value

OpenStudy (jazzyfa30):

ok

OpenStudy (rsadhvika):

f(x) = x2 +9

OpenStudy (rsadhvika):

MIMIMUM value = f(-b/2a) = f(0) = 0 +9 = 9

OpenStudy (rsadhvika):

if you dont get how to do this yet, dont wry.... u wil understand these in next problem im sure :)

OpenStudy (jazzyfa30):

yes

OpenStudy (rsadhvika):

so it will have a 'minimum value of 9'

OpenStudy (rsadhvika):

lets move on to next problem. hope we will dive in more deep in next prob ok :)

OpenStudy (jazzyfa30):

ok

OpenStudy (rsadhvika):

good :D